Can the converse of the Pythagorean theorem determine which triangle with the given three side lengths is a right triangle.
16,21,24
4,9,12
20,21,29
5,12,14

Found 2 solutions by mananth, Edwin McCravy:
Answer by mananth(16946)


16,21,24
choose the longest side. If it is a right triangle then 24 will be the hypotenuse.
Let's try
24 ^2 = 576 ----------(1)
16^2+21^ = 697----------(2)
(1) not equal to (2)
so it is not a right triangle
If they are equal then it is a right triangle



Answer by Edwin McCravy(20054)

16,21,24.
  
162 + 212 = 242
256 + 441 = 576
      697 = 576
       Nope!
4,9,12
42 + 92 = 122
16 + 81 = 144
     97 = 144
       Nope!
20,21,29
202 + 212 = 292
400 + 441 = 841
      841 = 841
       Yep!
5,12,14
52 + 122 = 142
25 + 144 = 196
     169 = 196
       Nope!

Edwin